# Now at ]Max-Planck-Institut für Gravitationsphysik (Albert-Einstein-Institut) and Leibniz Universität Hannover, 30167 Hannover, Germany Now at ]Max-Planck-Institut für Gravitationsphysik (Albert-Einstein-Institut) and Leibniz Universität Hannover, 30167 Hannover, Germany

### v1

@@ IV.1.4 Longitudinal Overlap
The longitudinal field overlap of the RC $\eta_{\hat{z}}$, expresses the relative reduction in the resonant enhancement factor experienced by a light field which is detuned from the cavity resonance by some frequency $\Delta\nu$, the difference between the frequency of the HPL and the nearest resonance of the RC. This is analogous to the square root of the ratio of the cavity transmissivity for a given $\Delta\nu$ versus the maximum transmissivity when $\Delta\nu=0$, or
$\eta_{\hat{z}}=\sqrt{\frac{T_{\rm RC}}{T_{\rm RC,max}}}.$ (61)
Therefore, $\eta_{\hat{z}}$ can be expressed as
@@
The expression above gives $\eta_{\hat{z}}$ in terms of a field overlap allowing it to take on some complex phase due to a non zero offset of the laser frequency from the cavity resonance. This will be useful to expressed the phase shift induced on the HPL signal as it transmits the RC when it is detuned from resonance. The approximation in Equation 62 can be expressed in terms of the measured parameters $\tau$ and $\Delta\nu$ as
@@
The error from using this approximation is on the order of 0.01% which is less than the other sources of error for the measurement.
Expressing $\eta_{\hat{z}}^{2}$ in terms of $\Delta\nu$ and $\nu_{0}$ (the cavities frequency half-width-half-maximum) gives
$\eta_{\hat{z}}^{2}\simeq\frac{1}{1+\left(\frac{\Delta\nu}{\nu_{0}}\right)^{2}}.$ (64)
In this equation it is clear that in the phrase ‘half-width-half-maximum’, the ‘half-maximum’ in this case refers to $\eta_{\hat{z}}^{2}$. With a mean value of $\tau$ of 6.85 ms for $\rm S_{{}_{\perp}}$ and 6.75 ms for $\rm S_{{}_{\parallel}}$, the mean value of the half-width-half-maximum of the RC ($\nu_{0}$) was 23.2 Hz for $\rm S_{{}_{\perp}}$ and 23.6 Hz for $\rm S_{{}_{\parallel}}$.
The phase change $\theta_{\hat{z}}$, experienced by the laser when it is transmitted through the cavity while off resonance can be found using the equation
@@
This phase change is important as it shows how a detuning of the HPL from the resonance of the cavity can distort the open shutter phase. This is one of the reasons why the calibration array is expressed in terms of a complex variable as it will be able to compensate for these phase changes when the closed shutter data are normalized.
@@
The initial FSR of the RC was used to establish the frequencies of the PLLs that control the frequency of the HPL relative to the frequency of the AL light transmitted by the RC. At 33 FSRs, this static frequency difference meant that the HPL frequency was tuned to be on resonance with the RC when the run was initially set up. However, as the length of the cavity drifted and the FSR changed, the HPL frequency was detuned from resonance and the longitudinal overlap $\eta_{\hat{z}}$, dropped below one. These effects also lead to a changes in $\theta_{\hat{z}}$, which was exhibited as a change in the phase of the open shutter measurements $C_{{\rm open},j}$.
The FSR of the RC was measured during maintenance periods by changing the frequency of PLL3 while the AL was stabilized to the cavity such that the LO was resonant at a frequency 45 FSRs away from the AL. The exact value of $\rm 45\cdot FSR$ was found at the frequency when the LO’s coupling to the cavity was optimized leading to a minimum in the amplitude of the beatnote between the two lasers in reflection of the cavity. With this technique it was possible to resolve the optimal beatnote frequency with a precision 1 Hz leading to a precision of the 0.02 Hz in the measured FSR. As the frequency spacing between AL and the HPL is 33 FSRs this means it is possible to measure the frequency detuning $\Delta\nu$ of the HPL with respect to the RC as a consequence of the changes in the FSR with a precision of 0.7 Hz.
@@
Figures 19(a) and 19(b) show $\Delta\nu$ for $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$. For $\rm S_{{}_{\perp}}$, the average absolute detuning of the HPL from the RC resonance (the mean value of $|\Delta\nu|$) estimated during the closed shutter periods was 5.4 Hz with a minimum detuning of 0 Hz, a maximum detuning of 14.5 Hz, and RMS deviations from the mean of 4.0 Hz. $\rm S_{{}_{\parallel}}$ showed an average absolute detuning of the HPL frequency of 8.9 Hz with a minimum detuning of 0.3 Hz, a maximum detuning of 28.4 Hz, an RMS of 6.4 Hz. The uncertainty in these measurements was 0.7 Hz as explained in the previous paragraph.
@@
Figures 19(c) and 19(d) show the projected HPL-RC longitudinal power overlap $\eta_{\hat{z}}^{2}$, from the HPL detuning data. The black vertical lines in Figure 19(d) show periods when the cavity length was adjusted to try to bring $\eta_{\hat{z}}^{2}$ back to one during $\rm S_{{}_{\parallel}}$. No such readjustments were performed on the length of the cavity during $\rm S_{{}_{\perp}}$. In these plots $\rm S_{{}_{\perp}}$ showed an estimated average value of $\eta_{\hat{z}}^{2}$ during the closed shutter periods of 92.8% with a minimum of 71.9%, a maximum value of 100%, and RMS fluctuations from the mean of 7.1%. For $\rm S_{{}_{\parallel}}$ the average value of $\eta_{\hat{z}}^{2}$ was 85.5%, while its minimum value was 40.6%, its maximum value was 100%, and its RMS was 14.4%. The uncertainty on these measurements were also primarily limited by the precision of the measurement of the RC FSR and was between 0.2% and 0.3%.
@@
The projections of $\theta_{\hat{z}}$ based on measurements of $\Delta\nu$, $\tau$, and $f_{0}$ are shown in Figures 19(e) and 19(f). Again, the black lines show periods where the cavity length was readjusted. For $\rm S_{{}_{\perp}}$, the mean absolute deviation of the longitudinal phase offset from zero estimated during the closed shutter periods was 0.22 rad and had a minimum value of 0 rad with a maximum value of 0.56 rad. The RMS fluctuations from the mean in this case were 0.16 rad. For $\rm S_{{}_{\parallel}}$ the average absolute deviation of $\theta_{\hat{z}}$ was 0.33 rad with a minimum of 0.01 rad, a maximum of 0.88 rad and an RMS of 0.22 rad. Like the previous calculations, the uncertainty in these measurements was mainly due to the uncertainty of the FSR measurement and was between 0.01 and 0.03 rad.
@@
It is interesting to compare the data shown in Figures 19(c), 19(d), 19(e), and 19(f), to their corresponding plots of the open shutter data $C_{{\rm open},j}$ measured at the science detectors in Figures 8(a), 8(b), and 9(a) and 9(b). Even though these two sets of plots were compiled from different sources of data, the relationship between $C_{{\rm open},j}$, and $\eta_{\hat{z}}$ and $\theta_{\hat{z}}$ is obvious. This demonstrates how the detuning of the HPL frequency from the RC was perhaps the most significant source of drifts in $C_{{\rm open},j}$.
Detuning the incident laser from the cavity resonance will also reduce the transmissivity and increase the reflectivity of the RC. Assuming perfect spatial overlap, the transmissivity $T_{\rm RC}$ as a function of the frequency detuning can be expressed as
$T_{\rm RC}=\eta_{\hat{z}}^{2}~T_{\rm RC,max},$ (67)
or, the product of the maximal transmissivity and the longitudinal overlap. During the closed-shutter periods of $\rm S_{{}_{\perp}}$, $T_{\rm RC}$ had a mean value of 3.3%, with RMS fluctuations about the mean of 0.2%, a maximum value of 3.6%, and a minimum value of 2.6%. The uncertainty of these measurements was 0.3%. For $\rm S_{{}_{\perp}}$, the closed-shutter periods gave a mean value of 3.0% for $T_{\rm RC}$, with RMS fluctuations of 0.5%, a maximum of 3.5%, and a minimum of 1.4%. The uncertainty was the same as the measurements during $\rm S_{{}_{\perp}}$ at 0.3%. In both runs the uncertainty was almost entirely driven by the uncertainty in the measurement of the transmissivity of the flat RC mirror mounted to the COB, RC2.
As Equations 55 and 56 showed, the reflectivity of the RC will also change as the input laser is detuned from resonance. Again assuming perfect spatial overlap, these equations can be expressed in terms of measured parameters as
$R_{{}_{\rm RC,NR}}\simeq\left|1-\frac{T_{{}_{\rm RC1}}f_{0}\tau}{1-i2\pi\tau\Delta\nu}\right|^{2}$ (68)
and
$R_{{}_{\rm RC,CH}}\simeq\left|1-\frac{T_{{}_{\rm RC2}}f_{0}\tau}{1-i2\pi\tau\Delta\nu}\right|^{2}.$ (69)
The quantity $R_{\rm RC,NR}$ is not used anywhere in this analysis as the HPL is incident on the cavity from CH and none of the lasers incident on the RC from NR experience the detuning from resonance. While $R_{\rm RC,CH}$ is used, the effect of the detuning is relatively muted as its estimated value during the closed shutter sections only varied from 92% to 94% over the duration of $\rm S_{{}_{\perp}}$ with an average value of 92% and RMS fluctuations from the mean of 0.6%. For $\rm S_{{}_{\parallel}}$ the RC reflectivity gave values between 91.6% to 96.6% with an average value of 93.0% and a and RMS of 1.2%. The uncertainty of these measurements was 4%, largely driven by the uncertainty in measurements of the transmissivity of RC2.
IV.2 Science Detection System
One of the most important metrics for evaluating the performance of the optical system is the spatial overlap between the HPL and the RC. This can be done by calibrating the results of the open shutter periods in terms of optical power and dividing by the expected power in transmission of the cavity. A number of parameters including the LO power in the spatial mode of the RC, the gain of the science detector, and the optical losses on the path from the RC to the photodetector must be measured to establish an expectation value for the open shutter power when the spatial coupling is perfect. Section IV.2.1 explains how $|C_{{\rm open},j}|^{2}$ can be expressed as a power in terms of these parameters, how these parameters are measured, and gives a summary of the data acquired during the first science campaign. Another essential component for assessing the power overlap between the HPL and the RC is the transmissivity of the optics mounted on the optical axis of the COB. The measurements of these optics are summarized in Section 5. Although errors in the polarization state of the HPL are not believed to be enough to be relevant to the result of the power overlap, an explanation of how these effects could couple into the power overlap is given in Section IV.2.3 as well as a rough estimate of their order based on measurements of the polarization state of the HPL. Finally, the measurements of the spatial power overlap at the science detector during the first science campaign are summarized in Section IV.2.4 and the methods used to optimize it are described in Section IV.2.5.
IV.2.1 Open Shutter Power
While $|C_{{\rm open},j}|^{2}$ gives a measure of the ratio between the signal power and the injected power, it is still in units of $\rm V^{2}/W$. For calculating the spatial overlap between the HPL and the RC it is useful to calculate the optical power of the HPL in transmission of the RC that is observed by the science detector during each open shutter period, as a function of other measured parameters. This power $P_{{\rm open},j}$, is
$P_{{\rm open},j}=\frac{P_{\rm i}|C_{{\rm open},j}|^{2}}{P_{\!{}_{\rm LO}}G_{\!{}_{\rm AC}}^{2}T_{\!{}_{\rm sci}}\eta_{\!{}_{xy,{\rm LO}}}^{2}}.$ (70)
@@ IV.2.1 Open Shutter Power
Here $P_{\!{}_{\rm LO}}$ is the LO (the laser used as the heterodyne local oscillator) power with units of W, $G_{\!{}_{\rm AC}}$ is the conversion gain of the science detection system including the full electronic chain at the first heterodyne demodulation frequency ($\sim$14 MHz) in V/W, $T_{\!{}_{\rm sci}}$ is the power transmissivity of the optical path from the RC to the science PD, and $\eta_{\!{}_{xy,{\rm LO}}}$ is the spatial overlap between the LO and the RC eigenmode.
@@
$P_{\!{}_{\rm LO}}$ is also an important parameter as it directly affects the amplitude of the heterodyne signal at the science detector. The change in the power during the runs is believed to be related to changes in the alignment of each laser to its corresponding optical fiber, as well as changes in $T_{\rm sci}$. For $\rm S_{{}_{\perp}}$ the mean value of $P_{\!{}_{\rm LO}}$ during the closed shutter runs was 310 µW with a range of 296 µW to 322 µW and RMS fluctuations from the mean of 10 µW. During $\rm S_{{}_{\parallel}}$ closed shutter periods $P_{\!{}_{\rm LO}}$ is estimated to have a mean value of 385 µW, with a minimum of 354 µW, a maximum of 408 µW, and an RMS of 15 µW. The uncertainty in the measurements of $P_{\!{}_{\rm LO}}$ during both runs was between 10 µW and 12 µW. Although both the uncertainty in the measurements of the DC voltage from the science detector, as well as the uncertainty in the measurement of the reflectivity of the RC, played a role in the overall uncertainty on $P_{\!{}_{\rm LO}}$, the dominant source of uncertainty in these measurements was the uncertainty in the measurements of the DC gain of the science detector electronic chain. This is discussed in more detail in the following section.
For the science detection system, it was important to measure both the DC conversion gain, as well as the gain of the system at the heterodyne frequency, $G_{\rm AC}$, to calibrate properly all of the parameters used for the measurement of $P_{{\rm open},j}$.
@@
The conversion gain of the science detector electronic chain at the heterodyne frequency $G_{\rm AC}$, was also measured under conditions meant to resemble the actual conditions of the science campaign. The LO and AL were both incident on the curved RC mirror in NR with PLL3 locked at its operating frequency. To make the measurements independent of the parameters of the RC the circulating field of the cavity was blocked between NR and CH and the AL power was reduced to 25.9 µW, a value near the mean of what was seen during the science campaign when the AL was stabilized to a resonance of the RC. A signal was sent to a fiber electro-optic modulator (EOM) to generate phase modulation sidebands on AL at the heterodyne frequency with roughly 0.01% of its carrier power. Therefore, like $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$, the light on the science PD had a large interference beatnote at the PLL3 frequency and a low amplitude signal at the heterodyne frequency. With this setup, the conversion gain of the full detection chain at $14.603$ MHz, $G_{\rm AC}$, could be measured. This was done using the ratio of the sideband to carrier beatnote amplitude measured at the PLL3 PD (see Figure 2) to estimate the relative power in the sideband. A previous measurement of the frequency-dependent gain of the PLL3 PD at these two frequencies was used to correct for the difference in the gain of the PD at 14.603 MHz and 54.95 MHz. The power of the LO was then measured at the science PD along with the amplitude of the signal at the heterodyne frequency. Thus the sideband power (with the sideband at the heterodyne frequency) could be calculated along with the conversion gain of the detection system, independent of the measurement of the DC conversion gain of the photodetector. These measurements resulted in a conversion gain of $2.9\pm 0.1$ V/mW. The primary uncertainty in these measurements was also related to the dependence of the gain on the position of the beam on the active area of the PD, as well as the calibration of the power meter.

### v2

@@ IV.2.1 Longitudinal Overlap
The longitudinal field overlap of the RC $\eta_{\hat{z}}$, expresses the relative reduction in the resonant enhancement factor experienced by a light field which is detuned from the cavity resonance by some frequency $\Delta\nu$, the difference between the frequency of the HPL and the nearest resonance of the RC. It can be expressed as
@@
This can also be expressed in terms of the measured parameters $\tau$ and $\Delta\nu$ as
@@
A laser that is detuned from the cavity resonance will also accumulate some additional phase offset during its transmission through the cavity. This phase offset $\theta_{\hat{z}}$, can be found using the equation
@@
This effect is important as it means a detuning of the HPL from the resonance of the cavity can distort the open shutter phase. This is one of the reasons why the calibration array is expressed in terms of a complex variable as it will be able to compensate for these phase changes when the closed shutter data are normalized.
@@
The initial FSR of the RC was used to establish the frequencies of the PLLs that control the frequency of the HPL relative to the frequency of the AL light transmitted by the RC. At 33 FSRs, this static frequency difference meant that the HPL frequency was tuned to be on resonance with the RC when the run was initially set up. However, as the length of the cavity drifted and the FSR changed, the HPL frequency was detuned from resonance and the longitudinal overlap $\eta_{\hat{z}}$, decreased. These effects also led to changes in $\theta_{\hat{z}}$, which was exhibited as a change in the phase of the open shutter measurements $C_{{\rm open}}$.
As Section IV.1.1 described, it was possible to measure the FSR with a precision a precision of the 0.02 Hz. The frequency detuning $\Delta\nu$ of the HPL with respect to the RC could therefore be measured with a precision of 0.7 Hz because the frequency spacing between AL and the HPL was 33 FSRs.
@@
Figures 19(a) and 19(b) show $\Delta\nu$ for $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$, as a result of the cavity FSR measurements. For $\rm S_{{}_{\perp}}$, the average absolute detuning of the HPL from the RC resonance (the mean value of $|\Delta\nu|$) estimated during the closed shutter periods was 5.4 Hz with a minimum detuning of 0 Hz, a maximum detuning of 14.5 Hz, and RMS deviations from the mean of 4.0 Hz. $\rm S_{{}_{\parallel}}$ showed an average absolute detuning of the HPL frequency of 8.9 Hz with a minimum detuning of 0.3 Hz, a maximum detuning of 28.4 Hz, an RMS of 6.4 Hz. The uncertainty in these measurements was 0.7 Hz as explained in the previous paragraph.
@@
Figures 19(c) and 19(d) show the projected HPL-RC longitudinal power overlap $\eta_{\hat{z}}^{2}$, from measurements of the cavity FSR and storage time. The black vertical lines in Figure 19(d) show periods when the cavity length was adjusted to try to bring $\eta_{\hat{z}}^{2}$ back to one during $\rm S_{{}_{\parallel}}$. No such readjustments were performed on the length of the cavity during $\rm S_{{}_{\perp}}$. In these plots $\rm S_{{}_{\perp}}$ showed an estimated average value of $\eta_{\hat{z}}^{2}$ during the closed shutter periods of 92.8% with a minimum of 71.9%, a maximum value of 100%, and RMS fluctuations from the mean of 7.1%. For $\rm S_{{}_{\parallel}}$ the average value of $\eta_{\hat{z}}^{2}$ was 85.5%, while its minimum value was 40.6%, its maximum value was 100%, and its RMS was 14.4%. The uncertainty on these measurements were also primarily limited by the precision of the measurement of the RC FSR and was between 0.2% and 0.3%.
@@
The projections of $\theta_{\hat{z}}$ based on measurements of the cavity FSR and storage time are shown in Figures 19(e) and 19(f). Again, the black lines show periods where the cavity length was readjusted. For $\rm S_{{}_{\perp}}$, the mean absolute deviation of the longitudinal phase offset from zero estimated during the closed shutter periods was 0.22 rad and had a minimum value of 0 rad with a maximum value of 0.56 rad. The RMS fluctuations from the mean in this case were 0.16 rad. For $\rm S_{{}_{\parallel}}$ the average absolute deviation of $\theta_{\hat{z}}$ was 0.33 rad with a minimum of 0.01 rad, a maximum of 0.88 rad and an RMS of 0.22 rad. Like the previous calculations, the uncertainty in these measurements was mainly due to the uncertainty of the FSR measurement and was between 0.01 and 0.03 rad.
@@
It is interesting to compare the data shown in Figures 19(c), 19(d), 19(e), and 19(f), to their corresponding plots of the open shutter data $C_{{\rm open}}$ measured at the science detectors in Figures 8(a), 8(b), and 9(a) and 9(b). Even though these two sets of plots were compiled from different sources of data, the relationship between $C_{{\rm open}}$, and $\eta_{\hat{z}}$ and $\theta_{\hat{z}}$ is obvious. This demonstrates how the detuning of the HPL frequency from the RC was perhaps the most significant source of drifts in $C_{{\rm open}}$.
IV.2.2 Polarization Overlap
The polarization overlap $\eta_{\rm pol}$, quantifies the mismatch between the ideal polarization state of the HPL for a given run ($\hat{x}$ for $\rm S_{{}_{\perp}}$ and $\hat{y}$ for $\rm S_{{}_{\parallel}}$) and the orientation of the magnetic field (defined to be in the $\hat{y}$ direction). The error in the orientation of the optical tables with respect to the magnetic fields is on the order of 100 µrad. The polarization state of the HPL for $\rm S_{{}_{\perp}}$ was tuned for maximum transmission through LT1 (which was rigidly mounted to the COB). For $\rm S_{{}_{\parallel}}$ an additional $\lambda/2$ wave-plate, whose orientation was precisely aligned, was also located on the COB in the path to LT1. The polarization state of the AL and LO are set by a rotatable half wave-plate in NR. The exact angle of the ‘roll’ of the COB with respect to the optical axis was not measured.
IV.2.3 Open Shutter Power Ratio
The ratio of HPL power leaving the cavity through $\rm RC1$ during the open shutter periods to the power injected to the production magnet string $P_{{\rm open}}/P_{{\rm i}}$, must be calculated to calibrate the field overlap measured at the science detector. While the open shutter calibration array from Section II.4 $|C_{{\rm open}}|^{2}$, gives a measure of the ratio between the signal power and the injected power, it is still expressed in terms of the ratio of the squared voltage measured by the heterodyne detection system to HPL power injected to the production magnet string (or units of $\rm V^{2}/W$). Nevertheless, $|C_{{\rm open}}|^{2}$ can be used with the expression
$\frac{P_{{\rm open}}}{P_{\rm i}}=\frac{|C_{{\rm open}}|^{2}}{P_{\!{}_{\rm LO}}G_{\!{}_{\rm AC}}^{2}T_{\!{}_{\rm sci}}\eta_{\!{}_{xy,{\rm LO}}}^{2}},$ (65)
@@ IV.2.3 Open Shutter Power Ratio
to calculate $P_{{\rm open}}/P_{{\rm i}}$ for a given open shutter period. Here $P_{\!{}_{\rm LO}}$ is the LO (the laser used as the heterodyne local oscillator), $G_{\!{}_{\rm AC}}$ is the conversion gain of the science detection system including the full electronic chain at the first heterodyne demodulation frequency ($\sim$14 MHz), $T_{\!{}_{\rm sci}}$ is the power transmissivity of the optical path from the RC to the science PD, and $\eta_{\!{}_{xy,{\rm LO}}}$ is the spatial overlap between the LO and the RC eigenmode.
@@
$P_{\!{}_{\rm LO}}$ is also an important parameter as it directly affects the amplitude of the heterodyne signal at the science detector. It was measured by calibrating the static voltage output by the science detector by the conversion gain and subtracting the power of the AL that was also present at this detector. The change in the power during the runs is believed to be related to changes in the alignment of each laser to its corresponding optical fiber, as well as changes in $T_{\rm sci}$. For $\rm S_{{}_{\perp}}$ the mean value of $P_{\!{}_{\rm LO}}$ during the closed shutter runs was 310 µW with a range of 296 µW to 322 µW and RMS fluctuations from the mean of 10 µW. During $\rm S_{{}_{\parallel}}$ closed shutter periods $P_{\!{}_{\rm LO}}$ is estimated to have a mean value of 385 µW, with a minimum of 354 µW, a maximum of 408 µW, and an RMS of 15 µW. The uncertainty in the measurements of $P_{\!{}_{\rm LO}}$ during both runs was between 10 µW and 12 µW. Although both the uncertainty in the measurements of the DC voltage from the science detector, as well as the uncertainty in the measurement of the reflectivity of the RC, played a role in the overall uncertainty on $P_{\!{}_{\rm LO}}$, the dominant source of uncertainty in these measurements was the uncertainty in the measurements of the DC gain of the science detector electronic chain. This is discussed in more detail in the following section.
@@
The conversion gain of the science detector electronic chain at the heterodyne frequency $G_{\rm AC}$, was also measured under conditions meant to resemble the actual conditions of the science campaign. The LO and AL were both incident on the curved RC mirror in NR with PLL3 locked at its operating frequency. To make the measurements independent of the parameters of the RC the circulating field of the cavity was blocked between NR and CH and the AL power was reduced to 25.9 µW, a value near the mean of what was seen during the science campaign when the AL was stabilized to a resonance of the RC. A signal was sent to a fiber electro-optic modulator (EOM) to generate phase modulation sidebands on AL at the heterodyne frequency with roughly 0.01% of its carrier power. Therefore, like $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$, the light on the science PD had a large interference beatnote at the PLL3 frequency and a low amplitude signal at the heterodyne frequency. With this setup, the conversion gain of the full detection chain at $14.603$ MHz, $G_{\rm AC}$, could be measured. This was done using the ratio of the sideband to carrier beatnote amplitude measured at the PLL3 PD (see Figure 2) to estimate the relative power in the sideband. A previous measurement of the frequency-dependent gain of the PLL3 PD at these two frequencies was used to correct for the difference in the gain of the PD at 14.603 MHz and 54.95 MHz. The power of the LO was then measured at the science PD along with the amplitude of the signal at the heterodyne frequency. Thus the sideband power (with the sideband at the heterodyne frequency) could be calculated along with the conversion gain of the detection system, independent of the measurement of the DC conversion gain of the photodetector. These measurements resulted in a conversion gain of $2.9\pm 0.1$ V/mW. This is slightly lower than the DC conversion gain as a result of the electronic transfer function of the photodetector. Like the measurements of the DC conversion gain, the primary uncertainty in these measurements was related to the dependence of the gain on the position of the beam on the active area of the PD, as well as the calibration of the power meter.